In renting an apartment with two other friends. Mike agreed to pay $100 security depost by himself. The 3 of them agreed to contribute equally toward the monthly rent. Mike's first check to the apt owner was for $425. What was the monthly rent for the apt?
Answer by mananth(16946)

In renting an apartment with two other friends. Mike agreed to pay $100 security depost by himself. The 3 of them agreed to contribute equally toward the monthly rent. Mike's first check to the apt owner was for $425. What was the monthly rent for the apt?
Let the monthly rent be x
each has to pay x/3 per month
Mike paid the deposit of $100
So he has to give a check of x/3 +100 which was 425
x/3 + 100 = 425
x/3 = 425-100
x/3 = 325
x= 975 rent for the month